# RxnCaption: Reformulating Reaction Diagram Parsing as Visual Prompt Guided Captioning

### v1

@@ 4.2 Main Results
We compared our RxnCaption-VL model with the baselines mentioned in § 4.1 on the RxnScribe-test and RxnCaption-11k-test, as shown in Table 4. It is evident that on both datasets, RxnCaption-VL achieved the best F1 scores, regardless of whether the HybridMatch or SoftMatch metric was used. Notably, on the more challenging RxnCaption-11k-test set, RxnCaption-VL outperformed the second-best model by 10.7 and 1.0 percentage points in F1 score for the HybridMatch and SoftMatch metrics, respectively, demonstrating the superior performance of RxnCaption-VL. Additionally, we observed that Gemini-2.5-pro exhibits excellent performance on the RxnDP task. When using the BIVP strategy, Gemini-2.5-Pro secured the highest F1 scores across almost all metrics on both test sets among the general VLMs. Specifically, on the RxnCaption-11k-test set, Gemini-2.5-pro achieved second place in F1 scores for both HardMatch and HybridMatch, even surpassing the specialized models RxnScribe and RxnIm.
@@ 4.2.1 BROS vs BIVP
As shown in Table 4, the BIVP strategy demonstrated a comprehensive advantage over the BROS strategy. BIVP consistently outperformed BROS across all general LVLMs. Among the specialized models, our RxnCaption-VL also significantly outperformed RxnScribe and RxnIm, which used the BROS strategy. We further compared the training of RxnCaption-VL using both BROS and BIVP strategies on the same training data. The results showed that the BIVP strategy achieved improvements of 9.3 and 4.6 percentage points in F1 score for the SoftMatch on the two test sets compared to the BROS strategy. These results convincingly demonstrate the effectiveness of our proposed BIVP strategy for the RxnDP task.
@@ 4.2.2 Real Data vs Synthetic Data
We further examined the differences between real paper data and synthetic data for the RxnDP task by comparing the performance of the RxnScribe, RxnCaption-VL, and RxnIm models. Despite the fact that the RxnIm training dataset (in terms of the number of images) is 38 times larger than that of RxnScribe and 4 times larger than that of RxnCaption-VL, its performance is noticeably inferior to these two models. Particularly on the RxnCaption-11k-test set, the SoftMatch F1 score for RxnIm was lower by 4.3 and 27.1 percentage points compared to RxnScribe and RxnCaption-VL, respectively. As discussed in § 3.2, the t-SNE visualization reveals a significant domain shift between RxnIm’s synthetic data and real data. These experimental conclusions further validate the limitations of current data synthesis methods.
4.2.3 Training RxnScribe with RxnCaption-11k
To validate the effectiveness of RxnCaption-11k, we retrained the RxnScribe model using the BROS strategy, while keeping the data configuration identical to that of RxnCaption-VL and following the official RxnScribe training scripts. This retrained model, referred to as RxnScribe_w/11k, was compared with the official open-source weights (RxnScribe_official), as shown in Table 4.2. The RxnScribe_w/11k showed significant improvement on the RxnCaption-11k-test set but a slight decrease on the original RxnScribe-test set. This decline is likely due to the limited capacity of Pix2Seq, which struggles to fully integrate the additional data. Moreover, constrained by the BROS strategy and model size, RxnScribe_w/11k still lags behind RxnCaption-VL.
@@ 4.3 Ablation Studies
Molecular Detector Rxn Extract RxnScribe-test RxnCaption-11k-test
Soft-F1 Hybrid-F1 Soft-F1 Hybrid-F1
MolYOLO Gemini 76.1 49.8 66.6 40.4
Ours 88.1 75.5 67.6 55.5
Ground Truth Bbox Gemini 87.4 53.0 78.8 47.4
Ours 89.7 76.7 71.3 58.5
MolYOLO Ground Truth Bbox 99.7 99.7 95.3 95.3
Table 5: Progressive ideal experiments for BIVP Strategy Error Analysis.
4.3.1 Error Analysis of BIVP Strategy
To identify error sources in the BIVP two-stage strategy, we conducted a progressive ideal experiment. During model inference, we sequentially replaced each stage’s predictions with ground truth annotations to quantify errors and assess the theoretical performance limit. The results are shown in Table 5.
In the first stage (molecular detection), we used the RxnCaption-VL training set strategy, inputting ground truth molecular bounding boxes as visual prompts. This resulted in only a slight improvement and even a minor performance drop on the RxnScribe-test set, indicating that molecular detection accuracy is not the primary bottleneck.
In the second stage (reaction extraction), we used bounding boxes predicted by MolYOLO to construct visual prompts, setting an ideal condition where the reaction is fully parsable if all molecular components are recalled (IoU >> 0.5). The results showed a significant improvement, with the F1 score on the RxnScribe test set nearing 100%. This confirms that reaction extraction is the main performance constraint in the current BIVP strategy, pointing to a key direction for future optimization.
@@ 4.3.2 Influence of Molecular Detector
LVLM Molecular Detector RxnScribe-test RxnCaption-11k-test
@@
Accurate molecular detection is essential for the BIVP strategy’s success. To validate our MolYOLO model, we substituted it with existing open-source detectors during the BIVP inference phase. Using RxnCaption-VL and Gemini models, we evaluated the impact of different detectors. The findings show that molecular detection capability greatly affects the BIVP strategy. Replacing MolYOLO with less effective detectors like YoDe or MolDetect resulted in a significant performance drop, as shown in Table 6.
@@ 5 Conclusion
This study presents the RxnCaption framework, enhancing LVLM performance in parsing chemical reaction diagrams by reformulating the task as image captioning through the “BBox and Index as Visual Prompt” (BIVP) strategy. Powered by our molecular structure detector MolYOLO, which achieves top accuracy on public benchmarks, BIVP provides a reliable visual prior. We developed the RxnCaption-11k dataset, a manually annotated collection of reaction diagrams from real chemical literature, significantly larger than existing datasets. Using this data and strategy, we trained the RxnCaption-VL model, setting new records in chemical reaction diagram parsing. Our methods, data, and models are expected to advance large-scale, high-precision extraction of reaction information from chemical literature, providing crucial technological support for the development of AI in chemistry.
@@ 6 Limitation
    Unresolved Challenges in Reaction Diagram Parsing: The task of parsing chemical reaction diagrams is not yet fully solved. Although models show generally high metrics on the RxnScribe dataset, the best-performing RxnCaption-VL model achieves an F1 score of less than 60% on our proposed RxnCaption-11k-test dataset. Continually improving model accuracy and performance is a key research focus. Additionally, effectively evaluating model uncertainty in challenging or currently unsolvable scenarios, enabling the model to abstain when necessary, is crucial for ensuring the reliability of output results and for facilitating the practical application of this technology.

### v2

@@ 4.2 Main Results
We compared our RxnCaption-VL model with the baselines mentioned in § 4.1 on the RxnScribe-test and RxnCaption-15k-test, as shown in Table 4. It is evident that on both datasets, our RxnCaption-VL (using BIVP) achieved the best F1 scores across both HybridMatch and SoftMatch metrics. Notably, on the more challenging RxnCaption-15k-test set, RxnCaption-VL outperformed the strongest competing model, Gemini-2.5-Pro (BIVP), by 19.4 percentage points in HybridMatch F1 and 3.8 percentage points in SoftMatch F1, demonstrating its superior performance. Additionally, we observed that Gemini-2.5-Pro exhibits excellent performance on the RxnDP task. When using the BIVP strategy, it secured the highest F1 scores among all general VLMs.
@@ 4.2.1 BROS vs BIVP
As shown in Table 4, the BIVP strategy demonstrated a comprehensive advantage over the BROS strategy. BIVP consistently outperformed BROS across all general LVLMs. We further compared the training of RxnCaption-VL using both BROS and BIVP strategies on the same training data. The results showed that on the RxnScribe-test set, the BIVP strategy improved the Hybrid-F1 by 3.0 percentage points and the Soft-F1 by a remarkable 10.0 percentage points compared to BROS. These results convincingly demonstrate the effectiveness of our proposed BIVP strategy for the RxnDP task.
@@ 4.2.2 Real Data vs Synthetic Data
We further examined the differences between real paper data and synthetic data for the RxnDP task by comparing the performance of the RxnScribe, RxnCaption-VL, and RxnIM models. Despite the fact that the RxnIM training dataset (in terms of the number of images) is 38 times larger than that of RxnScribe and 4 times larger than that of RxnCaption-VL, its performance is noticeably inferior to these two models. Particularly on the RxnCaption-15k-test set, the SoftMatch F1 score for RxnIM was lower by 5.4 percentage points compared to RxnScribe_official, and by a significant 29.9 percentage points compared to RxnCaption-VL. As discussed in § 3.2, the t-SNE visualization reveals a significant domain shift between RxnIM’s synthetic data and real data. These experimental conclusions further validate the limitations of current data synthesis methods.
4.2.3 Training RxnScribe with RxnCaption-15k
To validate the effectiveness of RxnCaption-15k, we retrained the RxnScribe model using the BROS strategy, while keeping the data configuration identical to that of RxnCaption-VL and following the official RxnScribe training scripts. This retrained model, referred to as RxnScribe_w/15k, was compared with the official open-source weights (RxnScribe_official), as shown in Table 4. The RxnScribe_w/15k showed significant improvement on the RxnCaption-15k-test set, with its Hybrid-F1 score increasing by 12.5 percentage points. However, on the original RxnScribe-test set, its Hybrid-F1 score showed a slight increase of 1.6 points, and the Soft-F1 score improved by 2.8 points. This indicates that while our dataset enhances performance on similar real-world data, the limited capacity of the Pix2Seq model may struggle to perfectly generalize across both data distributions. Moreover, constrained by the BROS strategy and model size, RxnScribe_w/1e5k still lags behind RxnCaption-VL.
@@ 4.3.1 Influence of Molecular Detector
LVLM Molecular Detector RxnScribe-test RxnCaption-15k-test
@@
Accurate molecular detection is essential for the BIVP strategy’s success. To validate our MolYOLO model, we substituted it with existing open-source detectors (YoDe, MolDetect) during the BIVP inference phase. As shown in Table 5, the choice of detector significantly impacts overall performance for both our model and the Gemini baseline.
For our model, switching from the weakest detector, YoDe, to our proposed MolYOLO boosts the Hybrid-F1 score by a remarkable 18.9 points on the RxnScribe-test set (from 53.3 to 72.2). This trend holds for the Gemini baseline as well, which sees a 15.1-point improvement under the same conditions. Even when compared to the strong MolDetect baseline, our MolYOLO provides a consistent performance lift across all metrics. These results clearly demonstrate that a high-quality molecular detector is a critical prerequisite for achieving state-of-the-art performance with the BIVP strategy, validating the contribution of our specialized MolYOLO detector.
Table 6: Error attribution analysis for the BIVP strategy. “MolYOLO” is the standard pipeline. “GT bbox + MolYOLO” refers to a setup where ground truth boxes are supplemented by MolYOLO’s non-overlapping predictions. RS=RxnScribe-test, RC-15k=RxnCaption-15k-test.
Detector Setup Extractor RS RC-15k
Hybrid-F1 Soft-F1 Hybrid-F1 Soft-F1
MolYOLO Gemini 49.8 76.1 40.4 66.2
Ours 72.2 86.2 59.8 70.4
GT bbox + MolYOLO Gemini 51.5 (+1.7) 82.2 (+6.1) 46.4 (+6.0) 73.5 (+7.3)
Ours 73.4 (+1.2) 87.6 (+1.4) 63.8 (+4.0) 76.3 (+5.9)
MolYOLO Ideal 99.7 99.7 95.3 95.3
4.3.2 Analysis of 2-Stage Strategy
To quantify the performance loss from each component of our BIVP strategy, we conducted an error attribution analysis, with results shown in Table 6. We designed a specific ideal experiment, “MolYOLO + GT bbox”, to analyze the errors. In this setup, we started with the complete set of ground truth bounding boxes, ensuring perfect recall. Then, we added any bounding box predicted by MolYOLO that does not have a significant overlap (IoU << 0.5) with any ground truth box. This setup effectively simulates a scenario with perfect recall but includes the noise from MolYOLO’s false positive detections and some free molecules.
By comparing this ideal setup with the standard MolYOLO pipeline, we can isolate and analyze the impact of the detector itself, where the improvement gained from the ideal detector is highlighted in red. As shown in the table, the gain is relatively small, especially on the simpler RxnScribe-test dataset. Even on the more challenging RxnCaption-15k-test set, the reaction extraction model remains the dominant bottleneck. Furthermore, when we assume an ideal reaction extraction model in Stage 2 (as shown in the last row of the table), the performance ceiling imposed by MolYOLO is only reduced by 0.3 and 4.7 points, respectively. This further reinforces our conclusion that enhancing the LVLM’s reasoning capability is the most critical direction for future improvements.
Besides, since BIVP adopts a two-stage design, a natural concern is whether it introduces additional computational overhead compared with end-to-end approaches. We test RxnCaption-VL on a single A100 (80GB). BROS requires 16.88 s/img, whereas BIVP achieves 9.65 s/img (0.10 s for MolYOLO + 9.55 s for the LVLM). The speedup comes from BIVP’s lightweight MolYOLO detector and the substantially shorter output format enabled by removing verbose bounding-box coordinates.
4.3.3 Analysis of Different Reaction Layout
Figure 5: Performance of RxnCaption-VL and Gemini on Different Reaction Layout.
Figure 5 highlights distinct patterns across reaction layout categories. Across reaction layout categories, we observe clear performance stratification, especially on RxnCaption-15k-test: single-line, multi-line and tree layouts show no substantial performance difference, while cyclic reactions exhibit drops. This progressive pattern highlights that structural complexity remains the primary bottleneck, suggesting substantial room for improvement in modeling complex reaction layouts.
Notably, despite the increased strictness of HybridMatch, our model maintains stable performance after training without experiencing degradation. This stability indicates that our model preserves structural consistency even under stringent matching criteria, further demonstrating its robustness to varying reaction complexities.
@@ 5 Conclusion
This study presents the RxnCaption framework, enhancing LVLM performance in parsing chemical reaction diagrams by reformulating the task as image captioning through the “BBox and Index as Visual Prompt” (BIVP) strategy. Powered by our molecular structure detector MolYOLO, which achieves top accuracy on public benchmarks, BIVP provides a reliable visual prior. We developed the RxnCaption-15k dataset, a manually annotated collection of reaction diagrams from real chemical literature, significantly larger than existing datasets. Using this data and strategy, we trained the RxnCaption-VL model, setting new records in chemical reaction diagram parsing. Our methods, data, and models are expected to advance large-scale, high-precision extraction of reaction information from chemical literature, providing crucial technological support for the development of AI in chemistry.
@@ 6 Limitation
    Unresolved Challenges in Reaction Diagram Parsing: The task of parsing chemical reaction diagrams is not yet fully solved. Although models show generally high metrics on the RxnScribe dataset, the best-performing RxnCaption-VL model achieves an F1 score of less than 60% on our proposed RxnCaption-15k-test dataset. Continually improving model accuracy and performance is a key research focus. Additionally, effectively evaluating model uncertainty in challenging or currently unsolvable scenarios, enabling the model to abstain when necessary, is crucial for ensuring the reliability of output results and for facilitating the practical application of this technology.
@@ 7 Acknowledgments
[…]
8 Related Work
8.1 Chemical Reaction Mining
Existing research on extracting chemical reactions from literature primarily employs text mining to identify entities such as reactants and products in patents [schneider2016big]. Tools like LeadMine perform entity recognition, while fingerprint-based methods enhance role assignment in noisy data [schneider2016s]. Rule-based and machine learning approaches, including Naive Bayes, are used for section identification and dataset construction [schneider2016big]. Recent initiatives like the ORD offer structured schemas for reaction data sharing [kearnes2021open]. Reaction Miner [zhong2023reaction] and automated methods using Transformer-based ChemBERT improve product identification and role labeling [guo2021automated]. ReactIE [zhong2023reactie] employs weak supervision with linguistic cues to enhance extraction. LLMs facilitate zero-shot NER, IUPAC-to-SMILES conversion, and atom mapping, adding 26% new reactions from patents [vangala2024suitability].
Chemical reaction diagram parsing has advanced recently. ReactionDataExtractor 2.0 [wilary2023reactiondataextractor] combines molecular and arrow detectors with text recognition, but its rule-based system struggles with generalization due to cumulative errors. RxnScribe [qian2023rxnscribe] redefines diagram parsing as a sequence generation task using Pix2Seq [chen2021pix2seq], creating a RxnDP dataset from real papers. RxnIm [RxnIm] introduces LVLMs, using synthetic data to address scarcity and employing a three-stage training strategy: 1) Pre-localization for molecular detection; 2) Full parsing with synthetic data; 3) Fine-tuning with real data. Despite high computational demands, it only marginally improves parsing accuracy over RxnScribe and struggles with out-of-distribution samples (as detailed in § 4.2 of the Main paper).
8.2 LVLM and Visual Prompt
Recently, Large Vision-Language Models (LVLMs) [hurst2024gpt, bai2023qwenvlversatilevisionlanguagemodel, chen2024internvl] built on Large Language Models (LLMs) [bai2023qwentechnicalreport, shao2022internnewlearningparadigm, touvron2023llama] have made significant strides. These models excel in visual perception [Liu_2024], visual question answering (VQA) [yu2024mmvetevaluatinglargemultimodal], and multimodal reasoning [yue2024mmmuprorobustmultidisciplinemultimodal, gao2025pm4benchparallelmultilingualmultimodal], with applications in medicine [li2023llavamedtraininglargelanguageandvision], autonomous driving [duan2024cityllavaefficientfinetuningvlms], remote sensing [pang2024vhmversatilehonestvision, muhtar2024lhrsbotempoweringremotesensing], and OCR [wei2024generalocrtheoryocr20].
Visual prompts [wu2024visual] provide input in visual forms (e.g., images, selections, markings) in LVLMs, enhancing fine-grained understanding and task adaptability. Unlike text prompts, visual prompts directly affect the input image, improving visual attention and reducing hallucinations. Set-of-Mark (SoM) [yang2023set] demonstrated the effectiveness of ”discrete markings” by overlaying symbols on images to guide GPT-4V in tasks like directional question-answering without fine-tuning, establishing a ”marking as prompt” zero-shot paradigm. Vip-llava [cai2024vip] allows models to understand arbitrary visual prompts, enabling region-specific question-answering without additional fine-tuning, and introduces the ViP-Bench for evaluation. Image-of-Thought (IoT) [zhou2024image] suggests models plan a multi-step visual-text reasoning chain, dynamically generating necessary masks or text markings using external tools (e.g., SAM [kirillov2023segment], OCR), achieving an automated “prompt as program” process.
9 Details for Pilot Study
9.1 Introduction of RxnCaption-15k
VQA Question Gemini-2.5-Pro GPT4o-2024-11-20 QwenVL-Max
How many reactions are in the image? 72.72 61.81 74.54
How many molecular structures are in the images? 66.36 25.45 56.36
Are there any cyclic reactions? 91.82 87.27 77.27
Are there any tree-structured reactions? 72.73 71.82 86.36
Average 75.91 61.59 73.63
Table 7: Zero-shot results of VQA task on RxnScribe-test-slct for General-purpose VLM
Prompt: BROS pattern
You are given an image containing one or more chemical reaction equations. Each equation has three parts: reactants, conditions, and products. Each part may include multiple objects, and each object is either a structure, text, identifier, or supplement. Please extract all objects and their bounding boxes, and return them in the following strict JSON format: a list, where each element represents a reaction with keys ’reactants’, ’conditions’, and ’products’. Each key maps to a list of objects, and each object has:
- category: one of ”structure”, ”text”, ”identifier”, or ”supplement”
- bbox: a list of four normalized values [x1, y1, x2, y2], representing the bounding box of the object relative to the image size. Each value must be between 0 and 1.
The coordinates must satisfy: x1 <x2, y1 <y2. (x1, y1) corresponds to the top-left corner, and (x2, y2) to the bottom-right corner of the object in the image.
  "reactants": [{"category": "structure",
        "bbox": [0.1, 0.2, 0.3, 0.4]}],
  "conditions": [{"category": "text",
        "bbox": [0.32, 0.21, 0.4, 0.25]}],
  "products": [{"category": "structure",
        "bbox": [0.45, 0.2, 0.6, 0.4]}]
Output only the JSON. Do not include any explanation or additional text.
<image>Now output your JSON format result:
Table 8: Prompt for BROS pattern
Prompt: Reaction Counting
You are given an image containing one or more chemical reaction equations. Your task is to count how many distinct reaction equations are present in the image. Output a JSON object with a single key ’reaction_count’ and an integer value indicating the number of reactions. For example: {”reaction_count”: 2}.
Do not include any explanation or additional text.
<image>Now output your JSON format result:
Table 9: Prompt template for counting reaction
Prompt: Molecular Structure Counting
You are given an image containing one or more chemical reaction equations. Your task is to count how many distinct molecule structures (i.e., chemical structure diagrams) are present in the image. Output a JSON object with a single key ’structure_count’ and an integer value indicating the number of molecular structures. For example: {”structure_count”: 4}.
Do not include any explanation or additional text.
<image>Now output your JSON format result:
Table 10: Prompt template for counting molecular structures
Model Hard Match Soft Match
P R F1 P R F1
Gemini-2.5-Pro 0.7 0.6 0.6 38.2 33.0 35.4
GPT4o-2024-11-20 0.4 0.3 0.3 0.4 0.3 0.3
QwenVL-Max 0 0 0 5.0 4.0 4.4
Table 11: Zero-shot results of RxnDP task on RxnScribe-test-slct for General-purpose VLM in BROS strategy
Model Hard Match Soft Match
P R F1 P R F1
Gemini-2.5-Pro 17.2 18.8 17.9 77.8 84.5 81.0
GPT4o-2024-11-20 9.9 9.7 9.8 59.7 55.7 57.6
QwenVL-Max 14.5 13.4 13.9 69.6 63.6 66.5
Table 12: Zero-shot results of RxnDP task on RxnScribe-test-slct for General-purpose VLM in BIVP strategy
Prompt: Cyclic Reaction Detection
You are given an image of a chemical reaction. Your task is to determine whether the image contains a cyclic (graph-style) chemical reaction diagram. Output a JSON object with a single key ’cyclic’ and a boolean value indicating whether the image contains a cyclic (graph-style) chemical reaction diagram. If the image contains a cyclic reaction, output: {”cyclic”: true}. Otherwise, output: {”cyclic”: false}.
Do not include any explanation or additional text.
<image>Now output your JSON format result:
Table 13: Prompt template for detecting cyclic chemical reactions
Prompt: Tree Reaction Detection
You are given an image of a chemical reaction. Your task is to determine whether the image contains a tree (tree-style) chemical reaction diagram. Output a JSON object with a single key ’tree’ and a boolean value indicating whether the image contains a tree (tree-style) chemical reaction diagram. If the image contains a tree reaction, output: {”tree”: true}. Otherwise, output: {”tree”: false}. Do not include any explanation or additional text.
<image> Now output your JSON format result:
Table 14: Prompt template for detecting tree-structured chemical reaction
Prompt: BIVP pattern
You are an expert in chemical image structure analysis. You will be given an image in which all ”molecular structures” have been boxed and numbered. Your task is to identify and reconstruct the chemical reaction(s) in the image based on these boxed structures. Please follow the rules below:
1. Each reaction must contain three fields: reactants, conditions, products (each is a List[Dict])
2. Each Dict must include:
- type: one of ”mol”, ”txt”, or ”idt”
- if type is ”mol”, second field is index (box ID);
- if type is ”txt” or ”idt”, second field is content (raw text)
3. Boxed molecular structures must be type ”mol”. Other elements must use ”txt” or ”idt”
4. Ignore decorations, arrows, and illustrations; only extract real elements
5. Clean up identifiers and retain only the core ID
6. If no reaction exists in the image, return an empty list []
7. Use arrow direction: tail = reactants, head = products
Example format:
    "reactants": [
    {"type": "mol", "index": 1},
    {"type": "txt", "content": "NaCl"}
    "conditions": [
    {"type": "mol", "index": 3},
    {"type": "txt", "content": "H2O, 25°C"}
    "products": [
    {"type": "mol", "index": 2},
    {"type": "idt", "content": "1a"}
Do not include any explanations, comments, or non-structured data.
<image> Please extract the list of reactions from the image. Return only the JSON reaction list without adding any other content.
Table 15: Prompt for BIVP pattern
We define the Reaction Diagram Parsing (RxnDP) task as $R=\mathcal{F}(I)$, where $I$ represents the input chemical reaction diagram, $\mathcal{F}$ is the RxnDP model, and $R=\{R_{i}\}$ is the set of all chemical reactions in the diagram.
Each reaction $R_{i}$ consists of three roles:
Existing RxnDP datasets have notable limitations: the RxnScribe dataset, derived from real papers, includes only 1378 samples, while the RxnIm dataset, though large, is artificially synthesized, limiting diversity and model generalization (see § 4.2 of the Main paper). To overcome these issues, we developed the RxnCaption-15k dataset—a large, high-quality, and diverse RxnDP dataset. And the training data of RxnCaption-15k is shown in Figure 6.
Figure 6: Visualization results of RxnCaption-15k’s training data.
9.2 Prompts for LVLM inference
The Prompt template for BROS pattern is shown in Table 11. The prompt template for VQA are shown in Table 9, 10, 13, 14. The prompt template for BIVP is shown in Table 12.
9.3 Metric for Pilot Study
HardMatch: A strict reaction-level matching strategy that requires all components of a reaction — including molecular structures and textual elements in reactants, products, and conditions — to be correctly matched. Each predicted component must have an IoU >0.5 with its ground truth counterpart. A reaction is counted as a true positive only if all associated components meet this criterion.
SoftMatch: A relaxed reaction-level matching strategy that ignores all text components and merges conditions into reactants during evaluation. A reaction is counted as a true positive only if all molecular structures in the reactants and products achieve IoU >0.5 with the ground truth. This strategy focuses solely on structural alignment.
9.4 Detailed Results for Pilot Study
The results of RxnDP task on RxnScribe-test-slct for General-purpose VLM in BROS strategy, VQA task, and RxnDP task in BIVP strategy are shown in Tables 11, 7, and 12, respectively.
10 Details of method
10.1 OCR details for Chemical Diagram Parsing
For the RxnCaption-15k dataset, each image contains annotated bounding boxes (bbox) that identify specific regions of interest within chemical diagrams. We extract the bounding box coordinates and crop the corresponding image regions, then apply a specialized OCR prompt(see Table 16) to each cropped region using Gemini-2.5-Pro.
Prompt: Gemini OCR
OCR_PROMPT = ”You are a master OCR assistant for chemistry, tasked with analyzing an image snippet from a chemical diagram. Your goal is to provide the most accurate text representation.
Follow these steps:
Step 1: Analyze the image content type.
  • Is it a graphical chemical structure? (e.g., contains bond lines, benzene rings, complex drawings).
  • Is it purely text? (e.g., can be typed on a keyboard).
Step 2: Provide output based on the content type.
  • If it is a graphical structure: You MUST return the special token [GRAPHICAL_STRUCTURE]. Do not attempt to OCR it.
  • If it is purely text: Transcribe the text with maximum precision. 1. Pay attention to details: Preserve case-sensitivity (e.g., Pd vs pd), all symbols (degree, /, -, +, (), and numbers. 2. Handle multi-line text: If text is on multiple lines, combine them into a single line separated by a space.
Examples:
  • Image shows a drawing of a benzene ring →Output: [GRAPHICAL_STRUCTURE]
  • Image shows a drawing of a molecule labeled ”1a” →Output: [GRAPHICAL_STRUCTURE]
  • Image shows the text (CH3)2N →Output: (CH3)2N
  • Image shows the text Pd/C, H2 →Output: Pd/C, H2
  • Image shows the text electrocatalyst →Output: electrocatalyst
  • Image shows text on two lines: NaH and THF →Output: NaH THF
Final Output Format: Return ONLY the transcribed text OR the [GRAPHICAL_STRUCTURE] token. Do not add explanations, markdown, or quotes.”
Output only the JSON. Do not include any explanation or additional text. <image>Now output your JSON format result:
Table 16: Prompt for Gemini OCR
10.2 Details of four layouts
The visualization of chemical reaction processes typically follows specific logical layouts to accurately convey their intrinsic chemical relationships. There are four primary types of layouts, as shown in Figure 7: 1) the Single-line Layout, which is characterized by the linear display of a series of continuous chemical transformations along a single path, sequentially arranging starting materials, intermediates, and products to form an unbroken reaction chain; 2) the Multiple-line Layout, which structurally breaks down a long reaction route into several consecutive linear segments arranged in sections; 3) the Tree Layout, which features multiple independent reaction pathways diverging from a common starting material or intermediate to different products, forming a one-to-many radial structure; and 4) the Cyclic Layout, which uses a closed-loop diagram to describe the process of a species undergoing a series of transformations and ultimately regenerating to its initial state.
Figure 7: Four layouts of Rxn Diagram.
11 Details of the experiments
11.1 Results Categorized into Four Layout Types
Figure 8: Results categorized into four layout types on RxnScribe-test and RxnCaption-15k-test.
We conducted a detailed analysis on four layout categories using four trained models from the main experiment table and one top-performing zero-shot VLM, Gemini-2.5-Pro. The results are shown in Figure 8. From the figures, we can observe that, when categorized by layout type, the difficulty of the RxnDP task follows the order: Single-line << Multi-line << Tree << Cyclic. Performance across all categories is generally slightly lower on RxnCaption-15k-test, which is due to our image crop being performed at the figure/table level, resulting in greater diversity and realism. In terms of individual model performance, our RxnCaption-VL demonstrates more balanced results across all categories, especially on RxnScribe-test, whereas Gemini-2.5-Pro exhibits a more pronounced imbalance in performance across different categories.
Besides, by comparing the pix2seq-based models (RxnScribe-official and RxnScribe_w/15k) with the VLM-trained models, we observe that traditional approaches maintain relatively stable performance on categories such as single-line and multi-line, but experience a significant drop when handling more complex images like tree and cyclic structures. This suggests that the VLM’s strong visual-language understanding capability can greatly benefit reaction diagram parsing tasks involving intricate graph structures. Moreover, by comparing the performance of RxnScribe-official and RxnScribe_w/15k across the two test sets, we observe that our proposed RxnCaption-15k not only yields substantial improvements on its corresponding test set, but also leads to noticeable gains on the RxnScribe-test. This demonstrates the strong generalizability of our constructed training dataset.
11.2 Detailed Results of Ablation Studies
For ablation experiments Error Analysis of BIVP Strategy and Influence of Molecular Detector in the main text, we present the detailed results in Table 17 and Table 18, which include the Precision, Recall, and F1 scores for each dataset and each evaluation metric.
Detector Setup Rxn Extractor RxnScribe-test RxnCaption-15k-test
Soft Match Hybrid Match Soft Match Hybrid Match
P R F1 P R F1 P R F1 P R F1
MolYOLO Gemini-2.5-pro 67.9 86.5 76.1 44.7 56.1 49.8 64.2 69.2 66.2 38.9 42.1 40.4
RxnCaption-VL 85.3 87.1 86.2 71.6 72.7 72.2 71.3 69.4 70.4 60.3 59.3 59.8
GT bbox + MolYOLO Recall Gemini-2.5-pro 74.5 91.7 82.2 45.9 58.7 51.5 70.6 76.6 73.5 44.7 48.3 46.4
RxnCaption-VL 87.2 88.1 87.6 73.3 73.5 73.4 75.9 76.7 76.3 63.8 63.9 63.8
MolYOLO Ground Truth 100.0 99.4 99.7 100.0 99.4 99.7 100.0 91.0 95.3 100.0 91.0 95.3
Table 17: Detailed results (Table 6 in full paper) of Error Attribution Analysis of BIVP Strategy in ablation studies.
LVLM Molecular Detector RxnScribe-test RxnCaption-15k-test
Soft Match Hybrid Match Soft Match Hybrid Match
P R F1 P R F1 P R F1 P R F1
Gemini-2.5-pro YoDe 49.4 60.7 54.5 30.1 41.0 34.7 58.9 59.9 59.4 35.0 37.1 36.0
MolDetect 64.3 82.6 72.3 42.5 54.9 47.9 59.2 62.7 60.9 36.7 38.9 37.8
MolYOLO 67.9 86.5 76.1 44.7 56.1 49.8 64.2 69.2 66.6 38.9 42.1 40.4
RxnCaption-VL YoDe 63.4 59.7 61.5 54.3 52.4 53.3 62.2 58.3 60.2 49.9 48.5 49.2
MolDetect 84.1 84.7 84.4 70.9 70.7 70.8 65.3 62.8 64.0 54.0 52.5 53.2
MolYOLO 85.3 87.1 86.2 71.6 72.7 72.2 71.3 69.4 70.4 60.3 59.3 59.8
Table 18: Detailed results (Table 5 in full paper) of Influence of Molecular Detector in ablation studies.